A Bibliography of Supercomputing ’2001

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

25 October 2010 Version 1.02

Abstract

This bibliography records articles presented at the Supercomputing ’2001 conference.

Title word cross-reference

8.61 [NKK01].

Acceleration [CB01]. Access [AFN^{+}01]. Achieving [BAN01]. Adaptable [USL01]. Adapting [MJD01]. Adaptive [BAN01]. Algebra [BL01]. Algorithm [Ada01, SV01]. Algorithms [MK01, NKV^{+}01]. AMR [WHK^{+}01]. Analysis [AKSS01, TMM^{+}01]. Application [ALPF01, BF01, CB01, KAH^{+}01, RN01]. Applications [AKSS01, FMM^{+}01, LTB01, WWHG01].

Applying [SCB01]. Approach [ÇA01]. Array [GPZ01]. ASCI [MJD01]. Atmospheric [LTD01]. Atomistic [NKV^{+}01]. Automatic [ALPF01].

Balancing [GS01, LTB01]. Based [RH01, FMM^{+}01, SMN01]. Benchmark [SBO01]. Beowulf [KT01]. Binary [LC01]. Blocking [JMCF01]. Body [BL01]. bypass [SWP01].

C [WA01]. Cactus [ADF^{+}01]. Calculation [MGKX01]. Calculations [WHK^{+}01]. Case [RN01]. CC [WA01]. CC-NUMA [WA01]. Cell [DH01]. Challenge [AFN^{+}01]. Circulation [LTD01]. Climate [AFN^{+}01]. Cluster [USL01]. Clusters [MOM^{+}01, RN01]. CO [ACM01]. Coarse [ÇA01, MK01]. Coarse-Grain [ÇA01]. Coast [LBL01]. Coastal [LBL01]. Codes [NPA01]. Communication [RRR01]. Communications [DH01]. Communities [TBAD^{+}01]. Component [FMM^{+}01]. Component-based [FMM^{+}01]. Compressing [LC01]. Computational [ALPF01]. Computations [MPR01]. Computer [NKK01]. Computers [BL01]. Computing [ACM01, ADF^{+}01, GHH^{+}01, SMN01, USL01]. Consistent [MGKX01]. copy [SWP01]. Core [LTD01]. Cosmology [BAN01]. Cost [USL01]. Cplant^{T M} [BF01]. Creating [TBAD^{+}01]. Cross [WWHG01]. Cross-Domain [WWHG01]. Cycle [RHK01].

Data [AFN^{+}01, AKSS01]. Decision [LC01]. Decomposition [ÇA01, ZS01]. Dedicated [MPR01]. Delivering [CB01]. Denver [ACM01]. Design [ALPF01, DP01, TNS^{+}01]. Detecting [FJ01]. Devices [MPR01, MOM^{+}01]. Diagram [LC01]. Direct [DP01]. Distributed [Ada01, ADF^{+}01, FJ01, LTB01, TMM^{+}01]. Domain [WWHG01, ZS01]. driven [SWP01]. Dual [LBL01]. Dual-Level [LBL01]. Dynamic [GS01, LTB01, WA01]. Dynamical [LTD01]. Dynamics [ALPF01, CUS^{+}01, NKK01].

Earth [KT01]. Effectiveness [USL01]. Efficient [ADF^{+}01, AKSS01, DP01, RHK01]. Effort [NPA01]. Element [LTD01]. EMP [SWP01]. Encoding [LC01]. Engine [PJ01]. Environment [FMM^{+}01]. Environments [ADF^{+}01]. Ethernet [SWP01]. Execution [ADF^{+}01, AKSS01]. Expansion [DH01, GPZ01]. Experiments [PBD^{+}01]. Exploration [PF01]. Extreme [BAN01].

Fast [LM01, TNS^{+}01]. fastDNAml [SHB^{+}01]. FeMn [CUS^{+}01]. FeMn/Co [CUS^{+}01]. File [WWHG01]. Files [LC01].

Fine [RHK01]. Fine-Grain [RHK01]. Fire [RSM01]. Fireplane [Cha01]. Fluid [ALPF01]. Fly [MGKX01]. FMPL [TNS^{+}01]. Formation [BAN01]. Framework [WHK^{+}01].

Gathering [TBAD^{+}01]. Gauss [Ada01]. General [LTD01]. Generating [MK01]. Generation [MOM^{+}01]. Ghost [DH01]. Gigabit [SWP01]. Global [PF01]. Globally [ZS01]. Globus [ADF^{+}01, MJD01]. GPFS [PTH^{+}01]. GrADS [PBD^{+}01]. Grain [ÇA01, RHK01]. Grande [SBO01]. Graphics [LM01, MOM^{+}01, PJ01]. Grid [AFN^{+}01, FMM^{+}01, LBL01, MJD01, PBD^{+}01, TBAD^{+}01]. Grids [MK01, PF01]. Groups [RRR01].

Hardware [LM01, MPR01, MOM^{+}01]. Heterogeneous [ADF^{+}01, MPR01]. High [ACM01, AFN^{+}01, RSM01, WWHG01]. High-Performance [AFN^{+}01, WWHG01]. HPC [CB01]. Hypergraph [ÇA01]. Hypergraph-Partitioning [ÇA01].

I/O [DP01, RN01, RHK01, TBAD^{+}01]. I/O-Efficient [DP01]. Implementation [PTH^{+}01, SHB^{+}01, TNS^{+}01]. Improve [WA01]. Improved [RSM01]. Improving [GPZ01]. Increased [CB01]. Increasing [JMCF01]. Indexing [PF01]. inference [SHB^{+}01]. Information [LC01]. Interactivity [PJ01]. Interconnect [Cha01]. Interfaces [CUS^{+}01]. Interval [GS01]. Interval-Newton [GS01]. Inverted [LC01]. IO [PTH^{+}01]. IO/GPFS [PTH^{+}01]. Irregular [GPZ01, NPA01, PGB^{+}01]. Itanium^{T M} [GHH^{+}01]. iterative [ZS01].

Java [SBO01]. JavaPSL [FJ01]. Jini

[SMN01]. Jini-based [SMN01].

Kerberos [MJD01].

Large [KAH^{+}01, KT01, PF01, WHK^{+}01]. Large-Scale [KAH^{+}01]. Launch [BF01]. LegionFS [WWHG01]. Level [LBL01]. Libraries [PBD^{+}01]. Library [RRR01, TNS^{+}01]. likelihood [SHB^{+}01]. line [SCB01]. Linear [BL01]. Linux [RN01]. Load [GS01, LTB01]. Locality [JMCF01, WA01]. Logic [CB01].

Magnetic [CUS^{+}01]. Management [RSM01]. Materials [NKV^{+}01]. Matrix [LM01]. maximum [SHB^{+}01]. MDM [NKK01]. Mechanics [BL01]. Mechanism [PGB^{+}01]. Memory [Ada01, RH01, TNS^{+}01]. Mesh [BAN01]. Message [GS01, SWP01, TNS^{+}01]. message-passing [TNS^{+}01]. Method [DH01]. Methods [MK01]. Minimal [NPA01]. Modeling [FJ01, KAH^{+}01, KT01, LBL01, RSM01]. Models [LTD01]. Molecular [NKK01]. MPI [PTH^{+}01]. MPI-IO [PTH^{+}01]. MPI-IO/GPFS [PTH^{+}01]. Multi [CUS^{+}01]. Multi-teraflops [CUS^{+}01]. Multiblock [LBL01]. Multigrid [Ada01, MK01]. Multilevel [MK01]. Multiple [AKSS01]. Multiplies [LM01]. Multiprocessors [WA01].

NaCl [NKK01]. Network [RHK01]. Networking [ACM01]. Networks [PGB^{+}01]. Newton [GS01]. Next [MOM^{+}01]. Next-Generation [MOM^{+}01]. NIC [SWP01]. NIC-driven [SWP01]. Non [ZS01]. Non-iterative [ZS01]. Non-overlapping [ZS01]. November [ACM01]. NUMA [WA01]. Numerical [BAN01, PBD^{+}01].

O [RN01, RHK01, TBAD^{+}01]. O-Efficient [DP01]. Ocean [LBL01]. On-line [SCB01]. On-the-Fly [MGKX01]. operations [TNS^{+}01]. Optimisation [FMM^{+}01]. Optimization [ALPF01, SV01]. Optimized [PTH^{+}01]. ORT [RRR01]. Orthogonal [RRR01]. OS-bypass [SWP01]. Overhead [RH01]. overlapping [ZS01].

Page [WA01]. Parabolic [ZS01]. Parallel [BL01, BF01, FJ01, GS01, GPZ01, MPR01, NPA01, PJ01, SCB01, SBO01, SHB^{+}01, TMM^{+}01, WHK^{+}01, ZS01]. Parallelism [LBL01]. Partial [GPZ01]. Partitioning [ÇA01]. Passing [GS01, SWP01, TNS^{+}01]. PC [MOM^{+}01]. PDE [DH01]. Co [CUS^{+}01]. GPFS [PTH^{+}01]. Performance [ACM01, AFN^{+}01, CB01, FJ01, KAH^{+}01, TMM^{+}01, WWHG01, SHB^{+}01]. phylogenetic [SHB^{+}01]. Placement [WA01]. Portal [KBG^{+}01, SMN01]. Potential [CB01]. Predictive [KAH^{+}01]. Primordial [BAN01]. Problem [AFN^{+}01, BL01]. Problems [DH01, FJ01, ZS01]. Processing [SV01]. Processor [GHH^{+}01, RRR01]. program [SHB^{+}01]. Programming [NPA01]. Programs [FJ01, TMM^{+}01]. Propagation [KT01]. Purpose [NKK01].

Quantum [BL01]. Query [AKSS01].

Real [PF01]. Real-time [PF01]. Reconfigurable [CB01]. Recursive [JMCF01]. Reducing [DH01]. Reductions [GPZ01]. Refinement [BAN01]. Regular [PF01]. Remote [AFN^{+}01, TNS^{+}01]. Removing [RH01]. Research [NKV^{+}01]. Resolution [BAN01, RSM01]. Resolving [BAN01]. Routing [PGB^{+}01].

s [NKK01]. SAMR [LTB01]. SAMRAI [WHK^{+}01]. SC2001 [ACM01]. Scalability [KAH^{+}01]. Scalable [BF01, LC01, NKV^{+}01, WWHG01]. ScaLAPACK [PBD^{+}01]. Scale [KAH^{+}01, KT01, WHK^{+}01]. SCALEA [TMM^{+}01]. Scaleable [PJ01]. Scaling [NPA01]. Scheduling [SCB01]. Science [KBG^{+}01]. Scientific [GHH^{+}01]. Search [SV01]. Secure [MJD01, WWHG01]. Seidel [Ada01]. Seismic [KT01]. Shared [RH01]. Signal [SV01]. Simulation [AFN^{+}01, NKV^{+}01, NKK01]. Skewing [JMCF01]. Smoothers [Ada01]. Software [RH01]. Software-Based [RH01]. Solution [BL01]. Solvers [DP01, ZS01]. Solving [DH01]. Sparse [BL01, DP01]. Special [NKK01]. Special-Purpose [NKK01]. Spectral [LTD01]. Spin [CUS^{+}01]. Stable [ZS01]. Star [BAN01]. Static [PF01]. Stealing [RHK01]. Steering [MGKX01]. Stochastic [SV01]. Strategies [GS01]. Structure [CUS^{+}01]. Structured [WHK^{+}01]. Studies [CUS^{+}01]. Study [RN01]. Suite [SBO01]. Sun [Cha01]. Supercomputing [MOM^{+}01]. Supporting [ADF^{+}01, WWHG01]. System [Cha01, SMN01, WWHG01]. Systems [LC01, LTB01].

Technologies [AFN^{+}01]. Temporal [JMCF01]. teraflops [CUS^{+}01]. Terascale [LTD01]. Tflop [NKK01]. Tflop/s [NKK01]. Three [BL01]. Three-Body [BL01]. Throttling [RHK01]. time [PF01]. Tomography [SCB01]. Tool [ALPF01]. top [PTH^{+}01]. Topology [PGB^{+}01]. TPC [WA01]. TPC-C [WA01]. Transactions [MGKX01]. Tuning [SCB01].

U.S. [LBL01]. Unstructured [Ada01]. Using [BL01, BAN01, CB01, GS01, GPZ01, TMM^{+}01, WHK^{+}01, LM01, MOM^{+}01].

Verification [MGKX01]. Very [PF01]. Visual [MOM^{+}01]. Volume [MOM^{+}01].

Wave [KT01]. Weather [RSM01]. Well [TBAD^{+}01]. West [LBL01]. within [FMM^{+}01]. Workloads [AKSS01].

XCAT [KBG^{+}01].

Zero [SWP01]. Zero-copy [SWP01].

References

ACM:2001:SHP

[ACM01] ACM, editor. SC2001: High Performance Networking and Computing. Denver, CO, November 10–16, 2001. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN 1-58113-293-X. LCCN ????

Adams:2001:DMU

[Ada01] Mark F. Adams. A distributed memory unstructured Gauss– Seidel algorithm for multigrid smoothers. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ???? URL http:/

                                                   /www.sc2001.org/papers/pap.
                                                   pap128.pdf.

Allen:2001:SEE

[ADF^{+}01] Gabrielle Allen, Thomas Dramlitsch, Ian Foster, Nicholas T. Karonis, Matei Ripeanu, Edward Seidel, and Brian Toonen. Supporting efficient execution in heterogeneous distributed computing environments with Cactus and Globus. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

                                                   URL http://www.sc2001.org/
                                                   papers/pap.pap301.pdf.

Allcock:2001:HPR

[AFN^{+}01] Bill Allcock, Ian Foster, Veronika Nefedova, Ann Chervenak, Ewa Deelman, Carl Kesselman, Jason Lee, Alex Sim, Arie Shoshani, Bob Drach, and Dean Williams. High-performance remote access to climate simulation data: A challenge problem for data grid technologies. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ???? URL http:/

/www.sc2001.org/papers/pap.
pap283.pdf.

Andrade:2001:EEM

[AKSS01] Henrique Andrade, Tahsin Kurc, Alan Sussman, and Joel Saltz. Efficient execution of multiple query workloads in data analysis applications. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap305.pdf.

Abramson:2001:ADO

[ALPF01] David Abramson, Andrew Lewis, Tom Peachey, and Clive Fletcher. An automatic design optimization tool and its application to computational fluid dynamics. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap202.pdf.

Bryan:2001:AER

[BAN01] Greg L. Bryan, Tom Abel, and Michael L. Norman. Achieving extreme resolution in numerical cosmology using adaptive mesh refinement: Resolving primordial star formation. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

                                        URL http://www.sc2001.org/
                                        papers/pap.pap176.pdf.

Brightwell:2001:SPA

[BF01] Ron Brightwell and Lee Ann Fisk. Scalable parallel application launch on Cplant^{T M}. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

                                        URL http://www.sc2001.org/
                                        papers/pap.pap263.pdf.

Baertschy:2001:STB

[BL01] Mark Baertschy and Xiaoye Li. Solution of a three-body problem in quantum mechanics using sparse linear algebra on parallel computers. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

                                        URL http://www.sc2001.org/
                                        papers/pap.pap285.pdf.

Catalyurek:2001:HPA

[ÇA01] Ümit V. Çatalyúrek and Cevdet Aykanat. A hypergraphpartitioning approach for coarsegrain decomposition. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ???? URL http:/

                                        /www.sc2001.org/papers/pap.
                                        pap210.pdf.

Caliga:2001:DAP

[CB01] David Caliga and David Peter Barker. Delivering acceleration: The potential for increased HPC application performance using reconfigurable logic. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap233.pdf.

Charlesworth:2001:SFS

[Cha01] Alan Charlesworth. The Sun Fireplane system interconnect. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap150.pdf.

Canning:2001:MTS

[CUS^{+}01] A. Canning, B. Ujfalussy, T. C.

Schulthess, X.-G. Zhang, W. A. Shelton, D. M. C. Nicholson, G. M. Stocks, Yang Wang, and T. Dirks. Multi-teraflops spin dynamics studies of the magnetic structure of FeMn/Co interfaces. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap274.pdf.

Ding:2001:GCE

[DH01] Chris Ding and Yun He. A ghost cell expansion method for reducing communications in solving PDE problems. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap291.pdf.

Dobrian:2001:DES

[DP01] Florin Dobrian and Alex Pothen.

The design of I/O-efficient sparse direct solvers. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ???? URL http:/

/www.sc2001.org/papers/pap.
pap261.pdf.

Fahringer:2001:MDP

[FJ01] Thomas Fahringer and Clóvis Seragiotto Júnior. Modeling and detecting performance problems for distributed and parallel programs with JavaPSL. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ???? URL http:/

                                        /www.sc2001.org/papers/pap.
                                        pap245.pdf.

Furmento:2001:OCB

[FMM^{+}01] Nathalie Furmento, Anthony Mayer, Stephen McGough, Steven Newhouse, Tony Field, and John Darlington. Optimisation of component-based applications within a grid environment. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

                                        URL http://www.sc2001.org/
                                        papers/pap.pap215.pdf.

Greer:2001:SCI

[GHH^{+}01] Bruce Greer, John Harrison, Greg Henry, Wei Li, and Peter Tang. Scientific computing on the Itanium^{T M} processor. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

                                        URL http://www.sc2001.org/
                                        papers/pap.pap266.pdf.

Gutierrez:2001:IPI

[GPZ01] Eladio Gutiérrez, Oscar Plata, and Emilio L. Zapata. Improving parallel irregular reductions using partial array expansion. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

                                        URL http://www.sc2001.org/
                                        papers/pap.pap260.pdf.

Gau:2001:PIN

[GS01] Chao-Yang Gau and Mark A.

Stadtherr. Parallel interval-Newton using message passing: Dynamic load balancing strategies. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap198.pdf.

Jin:2001:ITL

[JMCF01] Guohua Jin, John Mellor-Crummey, and Robert Fowler. Increasing temporal locality with skewing and recursive blocking. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap268.pdf.

Kerbyson:2001:PPS

[KAH^{+}01] D. J. Kerbyson, H. J. Alme, A. Hoisie, F. Petrini, H. J. Wasserman, and M. Gittings. Predictive performance and scalability modeling of a largescale application. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap255.pdf.

Krishnan:2001:XSP

[KBG^{+}01] Sriram Krishnan, Randall Bramley, Dennis Gannon, Madhusudhan Govindaraju, Jay Alameda, Richard Alkire, Timothy Drews, and Eric Webb. The XCAT science portal. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ???? URL http:/

/www.sc2001.org/papers/pap.
pap287.pdf.

Komatitsch:2001:MSW

[KT01] Dimitri Komatitsch and Jeroen Tromp. Modeling of seismic wave propagation at the scale of the Earth on a large Beowulf. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

                                        URL http://www.sc2001.org/
                                        papers/pap.pap267.pdf.

Luong:2001:COM

[LBL01] Phu Luong, Clay P. Breshears, and Le N. Ly. Coastal ocean modeling of the U.S. West Coast with multiblock grid and duallevel parallelism. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

                                        URL http://www.sc2001.org/
                                        papers/pap.pap162.pdf.

Lai:2001:CIF

[LC01] Chung-Hung Lai and Tien-Fu Chen. Compressing inverted files in scalable information systems by binary decision diagram encoding. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

                                        URL http://www.sc2001.org/
                                        papers/pap.pap338.pdf.

Larsen:2001:FMM

[LM01] E. Scott Larsen and David McAllister. Fast matrix multiplies using graphics hardware. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

                                        URL http://www.sc2001.org/
                                        papers/pap.pap313.pdf.

Lan:2001:DLB

[LTB01] Zhiling Lan, Valerie E. Taylor, and Greg Bryan. Dynamic load balancing of SAMR applications on distributed systems. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap250.pdf.

Loft:2001:TSE

[LTD01] Richard D. Loft, Stephen J.

Thomas, and John M. Dennis. Terascale spectral element dynamical core for atmospheric general circulation models. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap189.pdf.

Miller:2001:FCV

[MGKX01] David W. Miller, Jinhua Guo, Eileen Kraemer, and Yin Xiong. On-the-fly calculation and verification of consistent steering transactions. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ???? URL http:/

/www.sc2001.org/papers/pap.
pap165.pdf.

Moore:2001:AGK

[MJD01] Patrick C. Moore, Wilbur R.

Johnson, and Richard J. Detry. Adapting Globus and Kerberos for a secure ASCI grid. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap192.pdf.

Moulitsas:2001:MAG

[MK01] Irene Moulitsas and George Karypis. Multilevel algorithms for generating coarse grids for multigrid methods. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ???? URL http:/

/www.sc2001.org/papers/pap.
pap276.pdf.

Muraki:2001:NGV

[MOM^{+}01] Shigeru Muraki, Masato Ogata, Kwan-Liu Ma, Kenji Koshizuka, Kagenori Kajihara, Xuezhen Liu, Yasutada Nagano, and Kazuro Shimokawa. Next-generation visual supercomputing using PC clusters with volume graphics hardware devices. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap293.pdf.

Marongiu:2001:PDH

[MPR01] Alessandro Marongiu, Paolo Palazzari, and Vittorio Rosato. Parallel dedicated hardware devices for heterogeneous computations. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap212.pdf.

Narumi:2001:TMD

[NKK01] Tetsu Narumi, Atsushi Kawai, and Takahiro Koishi. An 8.61 Tflop/s molecular dynamics simulation for NaCl with a specialpurpose computer: MDM. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap205.pdf.

Nakano:2001:SAS

[NKV^{+}01] Aiichiro Nakano, Rajiv K. Kalia, Priya Vashishta, Timothy J. Campbell, Shuji Ogata, Fuyuki Shimojo, and Subhash Saini. Scalable atomistic simulation algorithms for materials research. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap113.pdf.

Nikolopoulos:2001:SIP

[NPA01] Dimitrios S. Nikolopoulos, Constantine D. Polychronopoulos, and Eduard Ayguadé. Scaling irregular parallel codes with minimal programming effort. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap184.pdf.

Petitet:2001:NLG

[PBD^{+}01] Antoine Petitet, Susan Blackford, Jack Dongarra, Brett Ellis, Graham Fagg, Kenneth Roche, and Sathish Vadhiyar. Numerical libraries and the Grid: The GrADS experiments with ScaLAPACK. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ???? URL http:/

/www.sc2001.org/papers/pap.
pap177.pdf.

Pascucci:2001:GSI

[PF01] Valerio Pascucci and Randall J.

Frank. Global static indexing for real-time exploration of very large regular grids. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

                                        URL http://www.sc2001.org/
                                        papers/pap.pap114.pdf.

Puente:2001:NRM

[PGB^{+}01] V. Puente, J. A. Gregorio, R. Beivide, F. Vallejo, and A. Ibañez. A new routing mechanism for networks with irregular topology. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

                                        URL http://www.sc2001.org/
                                        papers/pap.pap226.pdf.

Perrine:2001:PGI

[PJ01] Kenneth A. Perrine and Donald R. Jones. Parallel graphics and interactivity with the scaleable graphics engine. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

                                        URL http://www.sc2001.org/
                                        papers/pap.pap139.pdf.

Prost:2001:MIG

[PTH^{+}01] Jean-Pierre Prost, Richard Treumann, Richard Hedges, Bin Jia, and Alice Koniges. MPI-IO/GPFS, an optimized implementation of MPI-IO on top of GPFS. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ???? URL http:/

                                        /www.sc2001.org/papers/pap.
                                        pap186.pdf.

Radovic:2001:ROS

[RH01] Zoran Radović and Erik Hagersten. Removing the overhead from software-based shared memory. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ???? URL http:/

/www.sc2001.org/papers/pap.
pap314.pdf.

Ryu:2001:ENT

[RHK01] Kyung D. Ryu, Jeffrey K.

Hollingsworth, and Peter J. Keleher. Efficient network and I/O throttling for fine-grain cycle stealing. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ???? URL http:/

/www.sc2001.org/papers/pap.
pap120.pdf.

Ross:2001:CSA

[RN01] Robert Ross and Daniel Nurmi.

A case study in application I/O on Linux clusters. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap166.pdf.

Rauber:2001:OCL

[RRR01] Thomas Rauber, Robert Reilein, and Gudula Rünger. ORT — A communication library for orthogonal processor groups. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap207.pdf.

Roe:2001:HRW

[RSM01] Kevin Roe, Duane Stevens, and Carol McCord. High resolution weather modeling for improved fire management. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

                                        URL http://www.sc2001.org/
                                        papers/pap.pap286.pdf.

Smith:2001:PJG

[SBO01] L. A. Smith, J. M. Bull, and J. Obdrzálek. A parallel Java Grande benchmark suite. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

                                        URL http://www.sc2001.org/
                                        papers/pap.pap158.pdf.

Smallen:2001:AST

[SCB01] Shava Smallen, Henri Casanova, and Francine Berman. Applying scheduling and tuning to online parallel tomography. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

                                        URL http://www.sc2001.org/
                                        papers/pap.pap170.pdf.

Stewart:2001:PIP

[SHB^{+}01] Craig A. Stewart, David Hart, Donald K. Berry, Gary J. Olsen, Eric A. Wernert, and William Fischer. Parallel implementation and performance of fastDNAml — a program for maximum likelihood phylogenetic inference. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

                                        URL http://www.sc2001.org/
                                        papers/pap.pap191.pdf.

Suzumura:2001:JBC

[SMN01] Toyotaro Suzumura, Satoshi Matsuoka, and Hidemoto Nakada. A Jini-based computing portal system. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ???? URL http:/

/www.sc2001.org/papers/pap.
pap200.pdf.

Singer:2001:SSS

[SV01] Bryan Singer and Manuela Veloso. Stochastic search for signal processing algorithm optimization. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ???? URL http:/

/www.sc2001.org/papers/pap.
pap195.pdf.

Shivam:2001:EZC

[SWP01] Piyush Shivam, Pete Wyckoff, and Dhabaleswar Panda. EMP: Zero-copy OS-bypass NIC-driven Gigabit Ethernet message passing. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap315.pdf.

Thain:2001:GWC

[TBAD^{+}01] Douglas Thain, John Bent, Andrea Arpaci-Dusseau, Remzi Arpaci-Dusseau, and Miron Livny. Gathering at the well: Creating communities for Grid I/O. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap321.pdf.

Truong:2001:USP

[TMM^{+}01] Hong-Linh Truong, Thomas Fahringer Georg Madsen, Allen D. Malony, Hans Moritsch, and Sameer Shende. On using SCALEA for performance analysis of distributed and parallel programs. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ???? URL http:/

/www.sc2001.org/papers/pap.
pap243.pdf.

Tatebe:2001:DIF

[TNS^{+}01] Osamu Tatebe, Umpei Nagashima, Satoshi Sekiguchi, Hisayoshi Kitabayashi, and Yoshiyuki Hayashida. Design and implementation of FMPL, a fast message-passing library for remote memory operations. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap179.pdf.

Underwood:2001:CEA

[USL01] Keith D. Underwood, Ron R. Sass, and Walter B. Ligon, III. Cost effectiveness of an adaptable computing cluster. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap310.pdf.

Wilson:2001:DPP

[WA01] Kenneth M. Wilson and Bob B. Aglietti. Dynamic page placement to improve locality in CC-NUMA multiprocessors for TPC-C. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap241.pdf.

Wissink:2001:LSP

[WHK^{+}01] Andrew M. Wissink, Richard D. Hornung, Scott R. Kohn, Steve S. Smith, and Noah Elliott. Large scale parallel structured AMR calculations using the SAMRAI framework. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ???? URL http:/

/www.sc2001.org/papers/pap.
pap146.pdf.

White:2001:LSS

[WWHG01] Brian S. White, Michael Walker, Marty Humphrey, and Andrew S. Grimshaw. LegionFS: A secure and scalable file system supporting cross-domain highperformance applications. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ????

URL http://www.sc2001.org/
papers/pap.pap324.pdf.

Zhuang:2001:SGN

[ZS01] Yu Zhuang and Xian-He Sun.

Stable, globally non-iterative, non-overlapping domain decomposition parallel solvers for parabolic problems. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN ???? URL http:/

/www.sc2001.org/papers/pap.
pap190.pdf.